OpenStudy (anonymous):

Need help please There are three Gala apples and two Macintosh apples in a basket. You pick two apples from the basket. a) Develop a probability distribution X for the number of Gala apples selected

OpenStudy (anonymous):

There are 3 gala out of 5 apples, the 2 gala out of 4 apples. So to find the probability you must solve this 3/5x2/4=???

OpenStudy (anonymous):

i got 3/10? @gaos

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Awesome now I have a part b) if you could help me with that I would really appreciate it! It says what is the probability that only one apple picked will be a Gala apple? @gaos

OpenStudy (anonymous):

Let's suppose that the first is gala and the second is macntosh. So there are 3 gala out of 5 and 2 Macintosh out of 4 How much is 3/5x2/4=???

OpenStudy (anonymous):

3/10 @gaos

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But that was the answer to the first question? Im confused @gaos

OpenStudy (anonymous):

Yes, that happened because when you took one gala from the set it remains 2 out four. But in that set only have two Macintosh out of 4. Then, it turns out that the probability for both is the same

OpenStudy (anonymous):

Okay now in the way of writing it to show my work how would i got about that? Could i just write: 3gala/5 apples x 2 Macintosh/4 apples= 3/10

OpenStudy (anonymous):

@gaos

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

Awesome youre the best! Thank YOu @gaos

OpenStudy (anonymous):

Anytime sweetie